OpenStudy (anonymous):

Solve for x, 6 over 5 = 3x over x + 3.

OpenStudy (anonymous):

You cross multiply here, right?

OpenStudy (diyadiya):

\[\frac{6}{5}= \frac{3x}{x+3}\] Yes first cross multiply

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow 6(x + 3) = 3x(5) }\) \(\Large \color{purple}{\rightarrow 6x + 18 = 15x }\) \(\Large \color{purple}{\rightarrow 18 = 9x }\) \(\Large \color{purple}{\rightarrow x = 2 }\)

OpenStudy (anonymous):

15x = 6x + 18?

OpenStudy (diyadiya):

Right @fallingangel

OpenStudy (anonymous):

Thank you! ♥